OpenStudy (anonymous):

find the domain and range of the graphs: y=2^x y=x^2-4 y=x^2+4x+3 y=sqrt(x-2) y=sqrt(4-x^2) y=|x+3|-2

OpenStudy (anonymous):

all of them?

OpenStudy (anonymous):

first one domain, all real numbers. you can raise 2 to any power. range, positive numbers. 2 to any power is positive

OpenStudy (anonymous):

second one domain, all real numbers. you can square anything and add 4 to the result range, all numbers greater than or equal to -4. any numbers squared us at least 0, so once you subtract 4 the least it can be is -4

OpenStudy (anonymous):

thanks so much for your help so far (:

OpenStudy (anonymous):

third one domain is all real numbers. it is a polynomial range, the graph is a parabola that faces up. the vertex is (-2,-1) so minimum is -1 and range is \[[-1,\infty)\]

OpenStudy (anonymous):

fourth one \[\sqrt{x-2}\] you have to make sure that \[x-2\geq 0\] so \[x\geq 2\] is the domain. range is \[[0,\infty)\] since the radical sign indicates the positive root.

OpenStudy (anonymous):

fifth one \[\sqrt{4-x^2}\] is the only tricky one. RANGE is \[[0,\infty)\] for the same reason as last one.

OpenStudy (anonymous):

if you know that this is the graph of the upper half of a circle with center at the origin and radius 2, then you know that the range is \[[-2,2]\] if you do not know that, then you have to solve \[4-x^2\geq0\] which is not too hard but you have to be careful. it is a parabola that faces down with zeros at -2 and 2, so it will be positive between the zero (draw a mental picture) in any case the domain is \[-2\leq x\leq 2\]

OpenStudy (anonymous):

last one domian, all real numbers because you can take the absolute value of anything range \[[-2,\infty)\] because the absolute value is a least 0, so whole thing is at least -2

OpenStudy (anonymous):

hope reasoning is clear

OpenStudy (anonymous):

thannks so freaking much!! :D